OpenStudy (anonymous):

Find the inverse of f(x)=x/(x+1)

OpenStudy (anonymous):

ok, solve for x and then flip the x's and y's

OpenStudy (anonymous):

x=xy+y when i flip them i get: y=xy+x did i do my algebra wrong or what?

OpenStudy (anonymous):

youre on the right path, now factor out the x's

OpenStudy (anonymous):

x(y+1)=y

OpenStudy (anonymous):

wait, I just noticed something, you should have y=x-xy

OpenStudy (anonymous):

so now, after you factor out you have y=x(1-y) and now divide (1-y) on both sides and you get \[\frac{ y }{ 1-y }=x \]

OpenStudy (anonymous):

after that, simply switch the x and y's.

OpenStudy (anonymous):

why should i have y=x-xy?

OpenStudy (anonymous):

you start with